Two geological field teams are working in a remote area. A global positioning system (GPS) tracker at their base camp shows the location of the first team as 37 km away, 24° north of west, and the second team as 26 km away, 34° east of north. When the first team uses its GPS to check the position of the second team, what does it give for the second team's (a) distance from them and (b) direction, measured from due east?
Answer #1

Define a triangle with base camp labelled point A,

first team as point B, and second team as point C.

The three sides opposite these points are labelled

as a, b, and c respectively.

Initially considering A to be the origin, solve

for angle A within the triangle:

angle A = (90-24) + 34 = 100

By law of cosines:

a^2 = b^2 + c^2 - 2*b*c*cos(A)

a^2 = 26^2 + 37^2 - 2*26*37*cos(100)

a^2 = 2379.099

a = 48.78 km

Solve for angle B by law of sines:

sin(B)/b = sin(A)/a

sin(B)/26 = sin(100)/48.78

sin(B) = (sin(100)/48.78)*26

B = 31.67 degrees

Now considering point B to be the origin, we have

leg "c" of the triangle entering at 24 degrees south

of west (ie. -24 degrees) and leg "a" leaving at

31.67 degrees above that.

Thus, the *direction* of leg "a" is:

31.67 degrees - 24 degrees = 7.66 degrees (north of east).

Answer:

Question a) distance = 48.78 km

Question b) direction = 7.66 degrees (north of east)